# Characterizing the minimax rate of nonparametric regression under bounded convex constraints

### v3

@@ Abstract
We quantify the minimax rate for a nonparametric regression model over a convex function class $\mathcal{F}$ with bounded diameter. We obtain a minimax rate of ${\varepsilon^{\ast}}^{2}\wedge\mathrm{diam}(\mathcal{F})^{2}$ where
$\varepsilon^{\ast}=\sup\{\varepsilon>0:n\varepsilon^{2}\leq\log M_{\mathcal{F}}^{\operatorname{loc}}(\varepsilon,c)\},$
where $M_{\mathcal{F}}^{\operatorname{loc}}(\cdot,c)$ is the local metric entropy of $\mathcal{F}$ and our loss function is the squared population $L_{2}$ distance over our input space $\mathcal{X}$. In contrast to classical works on the topic (cf. yang1999information), our results do not require functions in $\mathcal{F}$ to be uniformly bounded in sup-norm. In addition, we prove that our estimator is adaptive to the true point, and to the best of our knowledge this is the first such estimator in this general setting. This work builds on the Gaussian sequence framework of neykov2022minimax using a similar algorithmic scheme to achieve the minimax rate. Our algorithmic rate also applies with sub-Gaussian noise. We illustrate the utility of this theory with examples including multivariate monotone functions, linear functionals over ellipsoids, and Lipschitz classes.
@@ 1 Introduction
This paper investigates minimax rates for nonparametric regression models under convex constraints. We build on the techniques of Neykov (2022), in which a lower bound is established using Fano’s inequality, and then provide a universal scheme that matches this lower bound. The minimax rate turns out to depend closely on the local geometry of the constraint set, in particular, in terms of its local packing numbers. We are able to handle cases where the function in $\mathcal{F}$ do not have a uniform sup-norm bound. [^1] While we can handle function spaces unbounded in sup norm constraints we are always assuming that the function space has a bounded $L_{2}(\mathbb{P}_{X})$ diameter. Our algorithmic upper bound also applies when the noise term is sub-Gaussian.
@@ 1.1 Problem Formulation and Notation
Let $\mathcal{F}\subset L_{2}(\mathbb{P}_{X})$ be a convex and complete class of real-valued functions on $\mathcal{X}\subseteq\mathbb{R}^{p}$. Suppose we observe $\{(X_{i},Y_{i})\}_{i=1}^{n}$ where for each $i\in[n]$, $X_{i}$ is independently drawn from some distribution $\mathbb{P}_{X}$ on $\mathcal{X}$ and $Y_{i}=\bar{f}(X_{i})+\xi_{i}\in\mathcal{Y}\subseteq\mathbb{R}$ where $\xi_{i}\sim\mathcal{N}(0,\sigma^{2})$ are i.i.d., $\sigma>0$ some fixed constant (which does not scale with the sample size), and $\bar{f}\in\mathcal{F}$ the true regression function. We introduce the notation $\vec{Y}=(Y_{1},\dots,Y_{n}),\vec{X}=(X_{1},\dots,X_{n})$. We also consider sub-Gaussian noise, where we say a random variable $Z$ is sub-Gaussian with mean $\mu$ and variance proxy $\tau^{2}$ if $\mathbb{E}[\exp(t(Z-\mu)]\leq\exp(t^{2}\tau^{2}/2)$ for all $t>0$ and write $Z\sim\mathrm{SG}(\mu,\tau^{2})$. In the sub-Gaussian setting, we will assume $\xi_{i}\sim\mathrm{SG}(0,\sigma^{2})$ are drawn i.i.d.
@@
The $L_{p}(\mathbb{P}_{X})$-norm can be defined analagously for $p\geq 2$. We define $d:=\mathrm{diam}(\mathcal{F}):=\sup\{\|f-g\|_{L_{2}(\mathbb{P}_{X})}:f,g\in\mathcal{F}\}$; we assume this quantity is finite throughout the paper. The empirical $L_{2}$-distance given input points $X_{1},\dots,X_{n}\sim\mathbb{P}_{X}$ is defined by
@@
As a short-hand, let $g(\vec{X})$ be the vector with $g(X_{i})$ in the $i$th component, and similarly for $f(\vec{X})$ and $\overline{f}(\vec{X}).$ Then $\|f-g\|_{L_{2}(\mathbb{P}_{n})}=n^{-1/2}\|f(\vec{X})-g(\vec{X})\|_{2}$, using the Euclidean $L_{2}$ norm for vectors in $\mathbb{R}^{n}$. Let $B(f,r,\|\cdot\|_{L_{2}(\mathbb{P}_{X})})=B(f,r)=\{g\in\mathcal{F}:\|f-g\|_{L_{2}(\mathbb{P}_{X})}\leq r\}$. Unless specified, we will always use the $L_{2}(\mathbb{P}_{X})$-norm for balls and abbreviate the notation as $B(f,r)$.
Our goal is to quantify the minimax rate
$\inf_{\widehat{f}}\sup_{\bar{f}\in\mathcal{F}}\mathbb{E}_{X,Y}\|\widehat{f}(\vec{X},\vec{Y})-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}$
and achieve it with a universal scheme depending only on the local geometry of $\mathcal{F}$. Here the infimum is over all measurable functions of the data $\widehat{f}\colon\mathcal{X}^{n}\times\mathcal{Y}^{n}\to\mathcal{F}$.
@@
Definition 1.1 (Local Entropy).
Let $M(r,\mathcal{F})$ be the largest cardinality of a $r$-packing set in the $L_{2}(\mathbb{P}_{X})$ norm of $\mathcal{F}$ (see Definition 5.4 Wainwright, 2019, e.g., for a definition of a packing set). Define for a constant $c>0$ the following:
$M_{\mathcal{F}}^{\operatorname{loc}}(\varepsilon,c)=\sup_{f\in\mathcal{F}}M(\varepsilon/c,B(f,\varepsilon)\cap\mathcal{F}).$
Then $\log M^{\operatorname{loc}}_{\mathcal{F}}(\varepsilon,c)$ is the local entropy of $\mathcal{F}$.
@@ 1.2 Related Works
In Yang and Barron (1999), the authors study minimax rates in terms of global entropies for density estimation, but apply their results to nonparametric regression. They obtain a similar result in their Theorem 6, where the minimax rate is $\varepsilon^{2}$ for $\varepsilon$ satisfying $\log M(\varepsilon,\mathcal{F})\asymp n\varepsilon^{2}$. However they require $\sup_{f\in\mathcal{F}}\|f\|_{\infty}<L$ for some $L>0$, while our results relax this assumption. As a consequence, we are able to derive minimax rates on linear functionals with unbounded sup norms. In addition, yang1999information impose an assumption on the parameter space which essentially forces the local and global metric entropies to coincide. We do not require this condition in our work.
@@
mendelson_2017_local_versus_global considers prediction problems for convex function classes, including ones with unbounded sup-norm. The minimax rate we consider can be viewed as a special case of this prediction setting. However, Mendelson bounds the error with high probability rather than in expectation. Moreover, mendelson_2017_local_versus_global requires an assumption that $\mathcal{F}$ is $L$-sub-Gaussian. This definition has two components. First, for each $p\geq 2$ and $f,g\in\mathcal{F}\cup\{0\}$, there must exist $L>0$ such that $\|f-g\|_{L_{p}(\mathbb{P}_{X})}\leq L\sqrt{p}\|f-g\|_{L_{2}(\mathbb{P}_{X})}$. Second, the centered canonical Gaussian process induced by $\mathcal{F}$ must be bounded. That is, to each $f\in\mathcal{F}$, we associate a Gaussian random variable $G_{f}$ with $\mathbb{E}[G_{f}]=0$ and $\mathbb{E}[G_{f}G_{g}]=\mathbb{E}[fg]$ for all $f,g\in\mathcal{F}$, and the assumption is that $\sup_{f}G_{f}$ is a bounded random variable almost surely. Our approach only uses the first assumption (and only for the unbounded sup-norm case) along with the diameter being bounded in the $L_{2}(\mathbb{P}_{X})$ norm. This assumption about the diameter may be easier to verify in practice compared to the boundedness of the Gaussian process. Another distinction with our approach is that mendelson_2017_local_versus_global does not match the lower and upper bounds on the error rate.
@@
Other works include agglomeration_yang_2004, which examines taking data-dependent convex combinations of $n^{\tau}$ for $\tau>0$ many candidate regression procedures to obtain a rate of order of the minimax rate plus a penalty term. agglomeration_yang_2004 requires the function to be bounded in sup-norm, but does relax assumptions on the distribution of the error term. yang_2014_high_dimensional_nonparam studies the nonparametric regression setting under assumptions of sparsity or additivity of the function class. They also require assumptions that the distribution on the input space $\mathcal{X}=[0,1]^{p}$ has a density lower bounded on a sub-hypercube (i.e., the data is not from a strictly lower dimensional subspace). The paper relies on global metric entropy conditions related to the minimax rate, as in yang1999information.
@@ 1.3 Organization
Our paper is organized as follows. In Section 2 we derive the minimax rate assuming that there is a uniform bound on the sup-norm of functions in $\mathcal{F}.$ Section 3 relaxes this assumption and only assumes the $L_{2}(\mathbb{P}_{X})$-diameter of $\mathcal{F}$ is finite. Section 4 refines the error-rate of the algorithm’s output to be adaptive to the true $\bar{f}$.
We then derive in Section 5 numerous existing and new results for particular examples of $\mathcal{F}$. One of the simplest examples is the set of linear functionals, i.e., when $K\subset\mathbb{R}^{p}$ is compact and convex, we consider functions of the form $X^{T}\beta$ and draw data $X_{1},\dots,X_{n}$ from an isotropic sub-Gaussian distribution. For $K$ being the $\ell_{1}$-ball and assuming the true $\beta^{\ast}$ is $s$-sparse, we use Sudakov minoration along with properties of the Gaussian width of tangent cones stated in chandrasekaran2012convex and amelunxen2014living to derive a minimax rate of order $\varepsilon\asymp\sqrt{s\log(d/s)/n}$, matching results from raskutti2011minimax. Following the ellipse example in neykov2022minimax, we again consider sub-Gaussian design and linear functionals over ellipsoids and recover a similar rate. We compare this to results from pathak2023noisy.
We then use existing results about the metric entropy from yang1999information and gao_Wellner_monotone and recover the well-known minimax rate for monotone functions. We similarly derive minimax rates for functions with smoothness assumptions, a well-studied setting (stone_1980).
2 Bounded Case
The bounded case assumes there is some $B_{\mathcal{F}}>0$ such that $\|f\|_{\infty}\leq B_{\mathcal{F}}$ for all $f\in\mathcal{F}$. In this case, we have $d\leq 2B_{\mathcal{F}}$ by the triangle inequality. We implicitly assume $\mathcal{F}$ is totally-bounded, but Remark 2.13 details how this is not a necessary assumption, including in the later unbounded and adaptive settings.
The structure of this section is as follows. First we derive a lower bound on the minimax rate using Fano’s inequality. In particular, we find that the minimax rate is at least $\varepsilon^{2}$ up to constants where $\varepsilon$ satisfies $\log M_{\mathcal{F}}^{\operatorname{loc}}(\varepsilon,c)>4(n\varepsilon^{2}/(2\sigma^{2})\vee\log 2)$.
To upper bound the minimax rate, we first construct an algorithm that outputs an estimator achieving a squared error rate of $\varepsilon^{2}$ when $\varepsilon$ satisfies a (discretized) condition of the form $n\varepsilon^{2}>2\log M_{\mathcal{F}}^{\operatorname{loc}}\left(\overline{C}\varepsilon,c\right)\vee\log 2$ for some constant $\overline{C}$.
To tie together these two bounds, we show that either the minimax rate is $d^{2}$, in which case our algorithm clearly outputs a function in $\mathcal{F}$ achieving this error rate, or it has an error rate ${\varepsilon^{\ast}}^{2}$ satisfying both the lower and upper bound conditions mentioned above. In fact, we establish the rate is ${\varepsilon^{\ast}}^{2}\wedge d^{2}$, where
$\varepsilon^{\ast}:=\sup\{\varepsilon:n\varepsilon^{2}\leq\log M_{\mathcal{F}}^{\operatorname{loc}}(\varepsilon,c)\}$ (2.1)
for some sufficiently large constant $c$.
2.1 Lower Bound
The lower bound follows from Fano’s Inequality:

### v5

# Characterizing the minimax rate of nonparametric regression under bounded star-shaped constraints

@@ Abstract
We quantify the minimax rate for a nonparametric regression model over a star-shaped function class $\mathcal{F}$ with bounded diameter. We obtain a minimax rate of ${\varepsilon^{\ast}}^{2}\wedge\mathrm{diam}(\mathcal{F})^{2}$ where
$\varepsilon^{\ast}=\sup\{\varepsilon\geq 0:n\varepsilon^{2}\leq\log\mathcal{M}_{\mathcal{F}}^{\operatorname{loc}}(\varepsilon,c)\},$
where $\log\mathcal{M}_{\mathcal{F}}^{\operatorname{loc}}(\cdot,c)$ is the local metric entropy of $\mathcal{F}$, $c$ is some absolute constant scaling down the entropy radius, and our loss function is the squared population $L_{2}$ distance over our input space $\mathcal{X}$. In contrast to classical works on the topic (cf. Yang and Barron, 1999), our results do not require functions in $\mathcal{F}$ to be uniformly bounded in sup-norm. In fact, we propose a condition that simultaneously generalizes boundedness in sup-norm and the so-called $L$-sub-Gaussian assumption that appears in the prior literature. In addition, we prove that our estimator is adaptive to the true point in the convex-constrained case, and to the best of our knowledge this is the first such estimator in this general setting. This work builds on the Gaussian sequence framework of Neykov (2022) using a similar algorithmic scheme to achieve the minimax rate. Our algorithmic rate also applies with sub-Gaussian noise. We illustrate the utility of this theory with examples including multivariate monotone functions, linear functionals over ellipsoids, and Lipschitz classes.
@@ 1 Introduction
This paper investigates minimax rates for nonparametric regression models under star-shaped constraints. We build on the techniques of Neykov (2022) and Prasadan and Neykov (2024) with a universal scheme whose error matches the well-known minimax lower bound due to Fano’s inequality. The minimax rate turns out to depend closely on the local geometry of the constraint set, in particular, in terms of its local packing numbers. We are able to handle cases where the functions in $\mathcal{F}$ do not have a uniform sup-norm bound. [^1] While we can handle function spaces unbounded in sup norm constraints we are always assuming that the function space has a bounded $L_{2}(\mathbb{P}_{X})$ diameter, defined in Section 1.1. Our algorithmic upper bound also applies when the noise term is sub-Gaussian.
@@ 1.1 Problem Formulation and Notation
Let $\mathcal{X}\subseteq\mathbb{R}^{p}$ and let $\mathbb{P}_{X}$ be some distribution on $\mathcal{X}$. Let $\mathcal{F}$ be a function class over $\mathcal{X}$. We observe $\{(X_{i},Y_{i})\}_{i=1}^{n}$ where for each $i\in\{1,\dots,n\}$, $X_{i}$ is independently drawn from $\mathbb{P}_{X}$ and $Y_{i}=\bar{f}(X_{i})+\xi_{i}\in\mathcal{Y}\subseteq\mathbb{R}$. Here $\bar{f}\in\mathcal{F}$ denotes the true regression function and the $\xi_{i}\sim\mathcal{N}(0,\sigma^{2})$ are i.i.d. noise terms for some fixed constant $\sigma>0$ (which does not scale with the sample size). We also consider sub-Gaussian noise, where we say a random variable $Z$ is sub-Gaussian with mean $\mu$ and variance proxy $\tau^{2}$ if $\mathbb{E}[\exp(t(Z-\mu)]\leq\exp(t^{2}\tau^{2}/2)$ for all $t>0$ and write $Z\sim\mathrm{SG}(\mu,\tau^{2})$. In the sub-Gaussian setting, we will assume $\xi_{i}\sim\mathrm{SG}(0,\sigma^{2})$ are drawn i.i.d.
@@
The $L_{p}(\mathbb{P}_{X})$-norm can be defined analogously for $p\geq 2$. We write $\mathcal{F}\subseteq L_{2}(\mathbb{P}_{X})$ if $\int|f(x)|^{2}\mathrm{d}\mathbb{P}_{X}(x)<\infty$ for all $f\in\mathcal{F}$. We define $d:=\mathrm{diam}(\mathcal{F}):=\sup\{\|f-g\|_{L_{2}(\mathbb{P}_{X})}:f,g\in\mathcal{F}\}$; we assume this quantity is finite throughout the paper. Let $B(f,r,\|\cdot\|_{L_{2}(\mathbb{P}_{X})})=B(f,r)=\{g\in\mathcal{F}:\|f-g\|_{L_{2}(\mathbb{P}_{X})}\leq r\}$. Unless stated otherwise, we will always use the $L_{2}(\mathbb{P}_{X})$-norm for balls and abbreviate the notation as $B(f,r)$. We write $B_{2}^{p}(\beta,r)$ for the Euclidean $\ell_{2}$-ball in $\mathbb{R}^{p}$ with radius $r$ and center $\beta\in\mathbb{R}^{p}$.
The empirical $L_{2}$-distance given input points $X_{1},\dots,X_{n}\sim\mathbb{P}_{X}$ is defined by
@@
We introduce the notation $\vec{Y}=(Y_{1},\dots,Y_{n}),\vec{X}=(X_{1},\dots,X_{n})$. We denote by $g(\vec{X})$ the vector with $g(X_{i})$ in the $i$th component, and similarly for $f(\vec{X})$ and $\bar{f}(\vec{X}).$ We can equivalently write our $L_{2}(\mathbb{P}_{n})$ distances as $\|f-g\|_{L_{2}(\mathbb{P}_{n})}=n^{-1/2}\|f(\vec{X})-g(\vec{X})\|_{2}$, using the Euclidean $L_{2}$ norm for vectors in $\mathbb{R}^{n}$.
We often abbreviate $\{1,\dots,n\}$ with $[n]$. We write $x\lesssim y$ (or $y\gtrsim x$) if there is some positive absolute constant $\alpha$ such that $x\leq\alpha y$. If $x\lesssim y$ and $y\lesssim x$, we write $x\asymp y$. We denote the $p\times p$ identity matrix by $\mathbb{I}_{p}$.
Definition 1.1.
We say a function class $\mathcal{F}$ is star-shaped if there exists an $f_{c}\in\mathcal{F}$, called the center of $\mathcal{F}$, such that for any $f\in\mathcal{F}$ and $t\in[0,1]$, $tf+(1-t)f_{c}\in\mathcal{F}$.
Henceforth we will assume $\mathcal{F}\subseteq L_{2}(\mathbb{P}_{X})$ is a star-shaped and complete class of real-valued functions on $\mathcal{X}\subseteq\mathbb{R}^{p}$. Our goal is to quantify the minimax rate
$\inf_{\hat{f}}\sup_{\bar{f}\in\mathcal{F}}\mathbb{E}_{\vec{X},\vec{Y}}\|\hat{f}(\vec{X},\vec{Y})-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}$
and achieve it with a universal scheme depending only on the local geometry of $\mathcal{F}$. Here the infimum is over all measurable functions of the data $\hat{f}\colon\mathcal{X}^{n}\times\mathcal{Y}^{n}\to\mathcal{F}$.
A crucial assumption we impose on $\mathcal{F}$ and $\mathbb{P}_{X}$ is the following: for some fixed constant $L>0$ and every integer $p\geq 2$ we have
$\|f-g\|_{L_{2p}(\mathbb{P}_{X})}^{2p}\leq\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}L^{p-1}p!/2.$ (1.1)
This will later yield us a Bernstein-like concentration inequality from which we can derive our upper bound. The requirement is implied by having a bound on the sup-norm of $\mathcal{F}$, or an $L$-sub-Gaussian condition (explained in the Related Works section), and is in fact strictly more general than either of these assumptions.
@@
Definition 1.2 (Local Metric Entropy).
We say a set $S$ is an $r$-packing set in the $L_{2}(\mathbb{P}_{X})$ norm for a function class $\mathcal{G}$ defined on $\mathcal{X}$ if $S\subseteq\mathcal{G}$ and for all distinct $f,g\in S$, $\|f-g\|_{L_{2}(\mathbb{P}_{X})}>r$. We denote by $\mathcal{M}(r,\mathcal{G})$ the maximal cardinality of an $r$-packing set in the $L_{2}(\mathbb{P}_{X})$ norm of $\mathcal{G}$. Given our function class $\mathcal{F}$, we define for a constant $c>1$ the following:
$\mathcal{M}_{\mathcal{F}}^{\operatorname{loc}}(\varepsilon,c)=\sup_{f\in\mathcal{F}}\mathcal{M}(\varepsilon/c,B(f,\varepsilon)\cap\mathcal{F}).$
Then $\log\mathcal{M}_{\mathcal{F}}^{\operatorname{loc}}(\varepsilon,c)$ is the local metric entropy of $\mathcal{F}$. We may omit the subscript $\mathcal{F}$ unless necessary.
The local metric entropy has been used in several existing works to obtain minimax rates for non-parametric regression, including Yang and Barron (1999, Section 7), Yang (2004), Wang et al. (2014, Section C.3.), and Mendelson (2017).
@@ 1.2 Related Works
In Yang and Barron (1999), the authors study minimax rates in terms of global entropies for density estimation, but apply their results to nonparametric regression. They obtain a similar result in their Theorem 6, where the minimax rate is $\varepsilon^{2}$ for $\varepsilon$ satisfying $\log\mathcal{M}(\varepsilon,\mathcal{F})\asymp n\varepsilon^{2}$. However, they require $\sup_{f\in\mathcal{F}}\|f\|_{\infty}<L$ for some $L>0$, while our results relax this assumption. As a consequence, we are able to derive minimax rates on linear functionals with unbounded sup norms. In addition, Yang and Barron (1999) impose an assumption on the parameter space which essentially forces the local and global metric entropies to coincide. We do not require this condition in our work.
@@
Mendelson (2017) considers prediction problems for convex function classes with bounded $L_{2}(\mathbb{P}_{X})$ diameter, including ones with unbounded sup-norm. The minimax rate we consider can be viewed as a special case of this prediction setting, although we have a much broader set of constraints. However, Mendelson bounds the error with high probability rather than in expectation. Moreover, Mendelson (2017, Definition 2.1) requires an assumption that $\mathcal{F}$ is $L$-sub-Gaussian. This definition has two components. First, for each $p\geq 2$ and $f,g\in\mathcal{F}\cup\{0\}$, there must exist $L>0$ such that $\|f-g\|_{L_{p}(\mathbb{P}_{X})}\leq L\sqrt{p}\|f-g\|_{L_{2}(\mathbb{P}_{X})}$. Second, the centered canonical Gaussian process induced by $\mathcal{F}$ must be bounded. That is, to each $f\in\mathcal{F}$, we associate a Gaussian random variable $G_{f}$ with $\mathbb{E}[G_{f}]=0$ and $\mathbb{E}[G_{f}G_{g}]=\mathbb{E}[fg]$ for all $f,g\in\mathcal{F}$, and the assumption is that $\sup_{f}G_{f}$ is a bounded random variable almost surely. We impose a strictly weaker condition, which is implied, for example, by an $L$-sub-Gaussian assumption or a bounded sup-norm, but neither of those two conditions are necessary. We do not impose an assumption about boundedness of the Gaussian process. Another distinction with our approach is that Mendelson (2017) does not precisely match the lower and upper bounds on the error rate. In addition, since every convex set is star-shaped, we generalize the shape constraints.
@@
Other works include Yang (2004), which examines taking data-dependent convex combinations of $n^{\tau}$ for $\tau>0$ many candidate regression procedures to obtain a rate of order of the minimax rate plus a penalty term. This work requires the function to be bounded in sup-norm, but relaxes assumptions on the distribution of the error term. Yang and Tokdar (2015) studies the nonparametric regression setting under assumptions of sparsity or additivity of the function class. They also require assumptions that the distribution on the input space $\mathcal{X}=[0,1]^{p}$ has a density lower bounded on a sub-hypercube (i.e., the data is not from a strictly lower dimensional subspace). The paper also uses global metric entropy conditions related to the minimax rate, as in Yang and Barron (1999). Zhao and Yang (2023) study a very general set of non-parametric problems allowing for more general noise distributions as well as the noise varying with $X$. Their results impose a bounded sup-norm assumption on the function class, however, in contrast to our work. Moreover, we are interested in scenarios where the richness condition in the authors’ Definition 3 may not hold, i.e., when the local and global entropies do not match for small $\varepsilon$.
The techniques in this paper are similar to Neykov (2022) in that we successively build finer local packings of the function class, resulting in an infinite tree of points that densely cover the set. We then traverse this tree using observed data. However, we must perform an additional pruning step to correct a gap found in the argument of Neykov (2022). The procedure is described in full detail in Prasadan and Neykov (2024), where the authors derive minimax rates for adversarially contaminated, sub-Gaussian mean location models with a star-shaped constraint. A drawback of these techniques is that the tree construction as well as traversal is not efficiently computable, at least in its stated form.
@@ 1.3 Organization
In Section 3, we give a summary of well-known results about lower bounds for the minimax rates. Section 3 establishes a matching upper bound in three steps: Section 3.1 introduces our local packing algorithm for $\mathcal{F}$; Section 3.2 derives useful properties of our key assumption (1.1) (with comparisons to $L$-sub-Gaussianity); Section 3.3 bounds our algorithm’s error and proves it attains the minimax rate. Then in Section 4, we prove that the error-rate of the algorithm’s output is adaptive to the true $\bar{f}$, albeit only when $\mathcal{F}$ is convex.
We then consider in Section 5 numerous applications of our theory. One of the simplest examples is the set of linear functionals: given a star-shaped and compact set $K\subset\mathbb{R}^{p}$, we consider functions of the form $X^{T}\beta$ for $\beta\in K$ and draw data $X_{1},\dots,X_{n}$ from an isotropic sub-Gaussian distribution. For $K$ being the $\ell_{1}$-ball and assuming the true $\beta^{\ast}$ is $s$-sparse, we use Sudakov minoration along with properties of the Gaussian width of tangent cones stated in Chandrasekaran et al. (2012) and Amelunxen et al. (2014) to derive the known minimax rate of order $\varepsilon\asymp\sqrt{s\log(d/s)/n}$, matching results from Raskutti et al. (2011) (see also Wang et al. (2014) for more general bounds). Following the ellipse example in Neykov (2022), we again consider sub-Gaussian design and linear functionals over ellipsoids and recover a novel rate. We compare this to the work of Pathak et al. (2024). We then derive known minimax rates for functions with smoothness assumptions, a well-studied setting (Stone, 1980, 1982). Lastly, we consider the example of monotone functions in general dimensions. We use existing results about the metric entropy from Yang and Barron (1999) and Gao and Wellner (2007) to recover the well-known minimax rate. We then analyze monotone functions that are piecewise constant on $k$ hyper-rectangles and show that our algorithm matches the adaptation rate of the least squares estimator established by Han et al. (2019).
2 Lower Bound
Our lower bound results all use existing and well-known information theoretic arguments, but we provide the details for completeness. The same argument using the local entropy appears in Yang and Barron (1999, page 1594), Yang (2004, Section 6), and Wang et al. (2014, Section C.3), for example. The upshot is that the minimax rate is at least $\varepsilon^{2}$ up to constants where $\varepsilon$ satisfies $\log\mathcal{M}_{\mathcal{F}}^{\operatorname{loc}}(\varepsilon,c)>4(n\varepsilon^{2}/(2\sigma^{2})\vee\log 2)$.
We start by restating Fano’s Inequality: